50. What is wrong with this proof? Theorem For every positive integer n = (n + /2. Basis Step: The formula is true for n =
Answer #1

Answer:

basic step fails in the proof
LHS = 1
RHS = (1 + 1/2)^2 / 2
= (3/2)^2 / 2
= 9/4 / 2
= 9/8
so, LHS != RHS for n=1